OpenStudy (anonymous):

What are all the values of k for which the integral from -3 to k of x^2dx=0?

OpenStudy (anonymous):

\[\int\limits_{-3}^{k} x^2dx=0\]

OpenStudy (danjs):

Integrate to begin

OpenStudy (danjs):

\[\frac{ x^3 }{ 3 }\] From -3 to k

OpenStudy (anonymous):

I dont know how to integrate so I dont know what to do next

OpenStudy (danjs):

You need to use the bounds and do F(k) - F(-3) = 0

OpenStudy (danjs):

where F(x) = x^3 / 3

OpenStudy (danjs):

\[0 = \frac{ 1 }{ 3 }k^3 - \frac{ (-3^3 }{ 3 }\]

OpenStudy (danjs):

The graph of x^2 from -3 to any value will give you a non-zero number, so the only k value that will result in zero area under the curve, would be the same (-3).

OpenStudy (danjs):

\[\int\limits_{-3}^{k}x^2dx = 0\] For k=-3